OpenStudy (anonymous):

Kelly tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," Kelly explains. Using complete sentences, demonstrate to Kelly how the equation is still solvable.

OpenStudy (anonymous):

@dan815 @Abhisar @abb0t @Ashleyisakitty @adrynicoleb @alphadxg @anonymous_user

OpenStudy (anonymous):

@bakonloverk @berrypicking @beccamarx19

OpenStudy (anonymous):

@Brialynn96 @brittany210

OpenStudy (anonymous):

@Compassionate @Cosmichaotic @ceggersjr @cram @compphysgeek @cookiemonster2412

OpenStudy (anonymous):

@Destinymasha @driftracer305 @D3xt3R @dhirensharma @dijahx3

OpenStudy (anonymous):

@esshotwired @Elsa213 @eric_d

OpenStudy (driftracer305):

well i guess it can be anything from 1/0 to..... 99/0 or whatever....... so since it will be in denominator the variable will be like so.... say for example:- \[\frac{ 5 }{ x } ........... x = 4/0......... \frac{ 5 }{ 4/0 }......... \frac{ 5*0 }{ 4 }.......\frac{ 0 }{ 4 }... = 0\] ................................

OpenStudy (anonymous):

But do i knwo that it can be solvable, how do i put that into words

OpenStudy (driftracer305):

say that u can take the denominator as a fraction as well........ so like i did 5 over (4 over zero)...........so zero goes up with 5 and 4 stays down.......that way u have 0 over 4 which is still solvable.............

OpenStudy (driftracer305):

@keishathelove101 ......... all good?

OpenStudy (anonymous):

yes